OpenStudy (jhulian):

please help.. find limit as thete approaches 0 of cos(6theta)tan(6theta)/theta

OpenStudy (irishboy123):

\[ \dfrac{cos(6\theta)tan(6\theta)}{\theta}\] that's \[ \dfrac{sin(6\theta)}{\theta}\] !!!!!!!!!!!!

OpenStudy (solomonzelman):

Just to add the last hit to what IrishBoy did, \(\color{#000000 }{ \displaystyle \lim_{\theta \to 0}\frac{\sin(6\theta)}{\theta}\quad \Longrightarrow \quad 6\cdot \lim_{\theta \to 0}\frac{\sin(6\theta)}{6\theta} }\)

OpenStudy (solomonzelman):

but, I think that without that it's pretty obvious.